# Privacy Preserving Federated Unsupervised Domain Adaptation with Application to Age Prediction from DNA Methylation Data

### v1

@@ 5.2 Confidence Score Computation
Once the predictions from the feature models are obtained by the aggregator, they are sent to the owner of the target domain. The owner then computes the confidence score for each feature based on the distribution predicted by the feature models. This step is performed entirely locally by the target domain owner.
@@
For a sample from the target domain, denoted as $\tilde{x}i$, and a given feature $f$, let $\tilde{x}{i,f}$ represent the value of feature $f$ in $\tilde{x}i$, and $\tilde{x}{i,\neg f}$ represent the values of all other features in $\tilde{x}i$. Given $\tilde{x}{i,\neg f}$, the feature model $g_{f}$ outputs a posterior distribution that describes the expected values of $\tilde{x}{i,f}$ according to the dependency structure of the source domain. For a GPR model, this posterior is a normal distribution, directly obtained from the closed-form solution shown in Eq. 3. To evaluate how well the observed value $\tilde{x}{i,f}$ fits the predicted distribution, we apply the confidence measure proposed by Jalali and Pfeifer [26]:
$c_{f}(\tilde{x}_{i})=2\Phi\left(\frac{|\tilde{x}_{i,f}-\mu_{g_{f}}(\tilde{x}_{i,\neg f})|}{\sigma_{g_{f}}(\tilde{x}_{i,\neg f})}\right)$ (7)
@@
where $\Phi$ denotes the cumulative distribution function of a standard normal distribution. This confidence represents the probability that a value as extreme as $\tilde{x}{i,f}$ or more, relative to $\mu{g_{f}}(\tilde{x}_{i,\neg f})$, occurs in the posterior distribution predicted by $g_{f}$. The overall confidence for feature $f$ in the target domain is defined as the average of $c_{f}(\tilde{x}_{i})$ across all target inputs:
$c_{f}=\frac{1}{m}\sum_{i=1}^{m}c_{f}(\tilde{x}_{i}).$ (8)
@@
For each feature, $c_{f}$ represents how well the source-domain dependencies of feature $f$ align with those in the target domain. Once the confidence scores for all features have been computed by the target domain owner, they are shared with the source domain clients.
5.3 Final Adaptive Model Training
@@
For the final prediction task on the target domain, we collaboratively train weighted elastic nets in a federated manner, ensuring the privacy of individual data owners is protected. The weights for these elastic net models are derived from the confidence scores computed by the target domain owner in the previous step. By using a weighted elastic net, we scale the contribution of each feature to the regularization term based on the feature weights, which are determined by the confidence scores from the target domain data.
@@
In this formulation, $w_{f}$ denotes the feature weights, and $\alpha$ controls the balance between the $L_{1}$ (lasso) and $L_{2}$ (ridge) penalties. The feature weights are calculated by the source domain owners using the confidence scores shared by the target domain owner. To compute these feature weights, we adopt the approach proposed by Handl et al. [6], where the feature weights $w_{f}$ are defined as:
$w_{f}=(1-c_{f})^{k}$ (11)
Here, $k$ is a hyper-parameter specified by the source domain owners, with $k>0$. This hyper-parameter determines how the confidence scores are transformed into feature weights. As $k$ increases, the model puts a progressively higher penalty on features with low confidence, while features with higher confidence are penalized less severely.
@@
As shown in Equations 9, 10, and 11, freda has three external parameters: the weighting parameter $k$, the proportion of $L_{1}$ and $L_{2}$ penalties in the weighted elastic net $\alpha$, and the regularization parameter $\lambda$. Following Handl et al. [6], we fix $\alpha=0.8$. For the weighting parameter $k$, we empirically evaluate the performance of our framework and adjust its value accordingly (Sections 2.4.3 and 2.5). The most challenging hyper-parameter to determine is the regularization parameter $\lambda$, for which we adopt the prior knowledge approach proposed in [6].
@@ 5.3.1 Optimal Lambda Prediction
The regularization parameter $\lambda$ determines the contribution of the regularization term $J(\beta)$ to the overall objective function of the model. In our case, since domain adaptation heavily relies on this regularization term, selecting an optimal value for $\lambda$ is critical. If $\lambda$ is too small, the penalty on the features becomes weak, and differences among feature weights may not significantly influence the model’s objective function. On the other hand, if $\lambda$ is too large, the redistribution of coefficients across features with different weights may overly dominate the objective function, preventing the model from learning meaningful representations from the source domain [6]. Handl et al. argue that, for any target domain $T$, the optimal value, $\lambda^{T}_{\text{opt}}$, depends on how much adaptation is needed for transfer between the source and target domain.
Traditionally, the optimal value for the regularization parameter $\lambda$ is determined through cross-validation. However, since our main objective is to perform unsupervised domain adaptation, using cross-validation to select the optimal $\lambda$ is counter-intuitive in this context, as cross-validation cannot be performed on the target domain due to the absence of labels and must instead be performed on the training (source) data. In light of this, we adopt the prior knowledge approach proposed in [6]. Here, the prior knowledge refers to side information available from other sources and not a prior distribution in the Bayesian sense.
The prior knowledge approach requires that side information about the similarities between the domains in the target domain is known by the owner of the target domain. In freda, this approach involves the owner of the target domain partitioning the indices $\{1,\cdots,l\}$ of all domains available in the target domain data, creating two subsets $X^{t}_{1}$ and $X^{t}_{2}$. Afterwards, for all $i\in X^{t}_{1}$, source domain owners collaboratively train separate weighted elastic nets in a federated manner, where during the training of each model the $\lambda$ value is selected by varying different possible values on a grid. Then, all these models trained with different values of $\lambda$ are sent to the owner of the target domain, where the target domain owner picks the model that leads to the lowest MAE on the target domains in $X^{t}_{1}$, assuming that the corresponding labels are available. Next, the owner of the target domain fits a simple linear model for the relationship between domain similarity and the optimal $\lambda$ values obtained from the weighted elastic nets provided by the source domain owners. Using this model, the owner of the target domain predicts the optimal $\lambda$ values for all domains in the remaining partition $X^{t}_{2}$. These predicted values are then sent to the source domain owners, where the source domain clients collaboratively train the final adaptive model in a federated manner using these values as the regularization term.
@@ 6 Security Analysis
In this section, we analyze the security guarantees of each phase of our framework under our assumptions.
6.1 Assumptions
@@ 6.2 Security of the Federated Feature Model Training
As described in Section 5.1, the training of feature models in our framework consists of two key steps: federated hyper-parameter optimization and federated GPR training.
@@ 6.2.1 Security of the Federated Hyper-parameter Optimization
This step is performed collaboratively between the source clients and the aggregator. The local data of the source clients are never exposed directly to either the aggregator or other data owners during this process. The only information exchanged during this step includes the optimal values of the prior variance on the coefficients $\sigma_{p}^{2}$ (as defined in Eq. 2) and the variance of the additive noise $\sigma_{n}^{2}$ (from the closed-form solution in Eq. 3). These values are computed locally by each source client for each feature.
The use of secure aggregation, implemented via zero-sum masking as formally proven in [14], ensures that the semi-honest aggregator cannot infer individual hyper-parameter values for any client. Instead, the aggregator only gains access to the aggregated sum of hyper-parameters for each feature, which contains no exploitable information about the local datasets of the source clients.
@@ 6.2.2 Security of the Federated GPR Training
This step is performed collaboratively between all participants, which include the target client, the source clients, and the aggregator. During this step, for each feature $f$ and the remaining data matrix $X_{\neg f}$ (which is the matrix after feature $f$ is excluded), both the target client and the source clients first mask their local data matrix $X_{\neg f}$ using a common mask matrix. This mask is generated by the clients using a shared seed, as described in Section 5.1.2. The aggregator then uses the masked matrices to compute the gram matrix, which is subsequently used to calculate the closed-form solution of the GPR model for feature $f$.
@@
The only information available to the aggregator is the resulting gram matrix, which is computed from the masked matrices. This process of masking, implemented in the FLAKE framework as formally proven in [13], ensures that the semi-honest aggregator cannot infer the original local data matrices $X_{\neg f}$ from either the target client or any source clients.
A critical step during this process is the computation of the term $K_{*}K^{-1}y$, which represents the mean of the GPR model’s closed-form solution. Here, $K_{*}$ is a part of the gram matrix corresponding to the output of the linear kernel between the target data and the source data, and $K^{-1}$ is the inverse of the output of the linear kernel between the source data itself. The term $y$ refers to the global feature vector corresponding to feature $f$ in the source dataset.
The source clients are not allowed to share their local feature vectors $X_{f}^{m}$ (where $X_{f}^{m}$ denotes the feature vector for source client $m$) directly with the aggregator because, if the full $y$ was revealed, the aggregator could reconstruct the complete local data of each source client after training across all features. To mitigate this, the aggregator computes the intermediate matrix $T=K_{*}K^{-1}$ from the gram matrix. This matrix $T$ is then partitioned into $M$ sub-matrices, where $M$ is the number of source clients, and the size of each sub-matrix $T_{m}$ corresponds to the number of local samples held by the source client $m$.
These sub-matrices are sent to their respective source clients. Each source client then multiplies their sub-matrix $T_{m}$ with their local feature vector $X_{f}^{m}$ to compute their contribution to the final mean vector. To preserve privacy, each source client masks their computed mean vector using a non-zero sum mask shared only with the target client. The masked mean vectors are sent back to the aggregator, who sums them to form the masked global mean vector $(K_{*}K^{-1}y)$. The aggregator then forwards the masked mean vector and the variance to the target client. The target client removes the mask on the mean vector and uses the resulting mean and variance vectors to compute confidence scores.
Aggregator’s Perspective.  The aggregator learns the number of local data samples held by each source client, as revealed by the size of each sub-matrix $T_{m}$. This information is required for the correct partitioning of $T$ but does not expose any details about the content of the local data samples. Additionally, the mean vectors received by the aggregator are masked, ensuring that no information about the source clients’ local feature vectors can be inferred.
Source Clients’ Perspective.  Even if a semi-honest source client accesses its portion of the intermediate product $K_{*}K^{-1}$, it cannot reconstruct or infer any information about other source clients’ local data. The masking and partitioning mechanisms prevent any cross-client data leakage.
Target Client’s Perspective.  The target client receives only the final closed-form GPR solution, comprising the mean and variance vectors. As formally proven in [13], this solution does not allow the target client to infer any details about the local data of any source client.
6.3 Security of the Confidence Score Computation
@@
The local data of the target client is never directly exposed to either the aggregator or the source clients during this process. The only information shared with the aggregator and source clients consists of the feature weights for each tissue in the target domain. These feature weight vectors are aggregate values, computed by the target client by averaging the feature weights of the samples belonging to each tissue within the target domain data. Consequently, it is not possible for the aggregator or the source clients to infer any information on specific samples in the target data.
Moreover, even if the aggregator or source clients had access to the raw, individual feature weight matrix for the entire target domain and, subsequently, the confidence score matrix for the entire target domain, reconstructing the target data would still be infeasible. This is because such reconstruction would require access to the mean of each feature model computed in the previous step. As explained in Section 6.2.2, only the target client possesses the mean vector for each feature model.
@@ 6.4 Security of the Optimal Lambda Prediction
Aggregator’s Perspective.  During the federated training process, the aggregator receives masked model updates from the source clients, which it aggregates to compute the coefficients of the global model for the next iteration. The use of secure aggregation, implemented via zero-sum masking as formally proven in [14], ensures that a semi-honest aggregator cannot infer individual updates from any source client. Instead, the aggregator only has access to the aggregated global model, which contains no exploitable information about the local datasets of the source clients.
@@
Target Client’s Perspective.  After the federated training process between the aggregator and the source clients, the target client receives only the coefficients of models trained using various regularization parameters. These model coefficients are aggregate values that do not reveal any specific information about the local data of any source client.
@@ 6.5 Security of the Final Adaptive Model Training
This step is performed collaboratively between the source clients and the aggregator. Throughout this process, the local data of the source clients is never directly exposed to either the aggregator or any other participant.
The federated training of the weighted elastic nets in this phase follows the same secure protocol as described in Section 6.4. The use of secure aggregation ensures that the semi-honest aggregator cannot infer individual updates from any source client, as only the aggregated global model is available to the aggregator. Similarly, source clients receive updated global model coefficients after each training round but do not gain access to any other source client’s local data or updates, preserving privacy across participants.
Since this phase involves the same secure mechanisms as the optimal lambda prediction phase, all privacy guarantees established in Section 6.4 apply here. Specifically, the use of masking and aggregation techniques ensures that neither the aggregator nor any source client can infer sensitive information about other participants’ local data.
@@ 7 Conclusion
In this article, we propose freda, a framework for privacy-preserving unsupervised domain adaptation suitable for high-dimensional data. Our framework utilizes federated Gaussian processes to learn the dependency structure between the source and the target domain data in a distributed environment, while using randomized encoding and secure aggregation to ensure the privacy of local data owners. The learned dependency structures between the domains are then used to collaboratively train a final adaptive model, where source and target domain owners train a weighted elastic net in a federated manner.
We tested the performance of our framework on real data, focusing on the problem of age prediction from DNA methylation data across various tissues. We compared the performance of our framework against the centralized state-of-the-art domain adaptation method wenda as well as the non-adaptive centralized method en-ls. Our experimental results show that our framework performs similarly to the centralized state-of-the-art method wenda, while effectively addressing the distribution shift between the source and target domains, even in a distributed setting where the source domain data is spread across up to $8$ different parties.
This article demonstrates that unsupervised domain adaptation for high-dimensional data is not only feasible in a distributed setting but can also be achieved while protecting sensitive information through privacy-enhancing technologies. Unlike previous centralized approaches, our proposed framework is the first to address the challenges of privacy-preserving collaboration in high-dimensional domain adaptation. This is particularly relevant in fields such as computational biology and medicine, where institutions frequently work with sensitive data and face increasing demands for secure cross-institutional collaboration.

### v2

@@ 5.2 Feature Weight Computation
This step is performed entirely by the target client. After obtaining the predictions from the feature models, the target client computes the confidence score for each feature based on the distribution predicted by the feature models. These confidence scores are then used to determine the weight of each feature.
@@
For a sample from the target domain, denoted as $\tilde{x}_{m}$, and a given feature $f$, let $\tilde{x}_{m,f}$ represent the value of feature $f$ in $\tilde{x}_{m}$, and $\tilde{x}_{m,\neg f}$ represent the values of all other features in $\tilde{x}_{m}$. Given $\tilde{x}_{m,\neg f}$, the feature model $g_{f}$ outputs a posterior distribution that describes the expected value of $\tilde{x}_{m,f}$ according to the dependency structure learned from the source domain. For a GPR model, this posterior is a normal distribution, which is directly obtained from the closed-form solution shown in Eq. 2 and collaboratively computed in the previous phase.
To evaluate how well the observed value $\tilde{x}_{m,f}$ fits the predicted distribution, we apply the confidence measure proposed by Jalali and Pfeifer [39]:
$c_{f}(\tilde{x}_{m})=2\Phi\left(\frac{|\tilde{x}_{m,f}-\mu_{g_{f}}(\tilde{x}_{m,\neg f})|}{\sigma_{g_{f}}(\tilde{x}_{m,\neg f})}\right)$ (6)
@@
where $\Phi$ denotes the cumulative distribution function of a standard normal distribution. The terms $\mu_{g_{f}}$ and $\sigma_{g_{f}}$ correspond to the mean and standard deviation of the normal distribution predicted by the GPR model $g_{f}$, respectively. This confidence score represents the probability that a value as extreme as $\tilde{x}_{m,f}$, or more, relative to the predicted mean $\mu_{g_{f}}(\tilde{x}_{m,\neg f})$, occurs within the posterior distribution predicted by $g_{f}$.
The overall confidence for feature $f$ in the target domain is then defined as the average of $c_{f}(\tilde{x}_{m})$ across all target domain samples:
$c_{f}=\frac{1}{n_{t}}\sum_{i=1}^{n_{t}}c_{f}(\tilde{x}_{m})$ (7)
@@
Where $n_{t}$ represents the total number of samples in the target domain. For each feature, $c_{f}$ quantifies how well the source-domain dependencies for feature $f$ align with those in the target domain. Once the confidence scores for all features have been computed, the target client then computes the weight of feature $f$ as follows:
$w_{f}=(1-c_{f})^{k}$ (8)
Here, $k$ is a hyper-parameter specified by the target client, with $k>0$. This hyper-parameter determines how the confidence scores are transformed into feature weights. As $k$ increases, progressively higher penalties are applied to features with low confidence, while features with higher confidence are penalized less severely.
In our experiments, we empirically evaluate the performance of our framework with respect to the weighting parameter $k$ and adjust its value accordingly (Sections 2.4.3 and 2.5).
5.3 Federated Weighted Elastic Net Training
@@
The remaining phases of our framework involves collaboratively training weighted elastic nets in a federated manner, to preserve the privacy of individual source clients. By using a weighted elastic net, source clients scale the contribution of each feature in their local data to the regularization term based on the feature weights computed by the target client in the previous step.
@@
As shown in Equations 9, 10, freda has two external parameters: the proportion of $L_{1}$ and $L_{2}$ penalties in the weighted elastic net $\alpha$, and the regularization parameter $\lambda$. Following Handl et al. [20], we fix $\alpha=0.8$. The most challenging hyper-parameter to determine is the regularization parameter $\lambda$, for which we adopt the prior knowledge approach proposed in [20].
@@ 5.3.1 Optimal Lambda Prediction
The regularization parameter $\lambda$ determines the contribution of the regularization term $J(\beta)$ to the overall objective function of the model. In our case, since domain adaptation heavily relies on this regularization term, selecting an optimal value for $\lambda$ is critical. If $\lambda$ is too small, the penalty on the features becomes weak, and differences among feature weights may not significantly influence the model’s objective function. On the other hand, if $\lambda$ is too large, the redistribution of coefficients across features with different weights may overly dominate the objective function, preventing the model from learning meaningful representations from the source domain [20].
Traditionally, the optimal value for the regularization parameter $\lambda$ is determined through cross-validation. However, since our main objective is to perform unsupervised domain adaptation, using cross-validation to select the optimal $\lambda$ is counter-intuitive in this context, as cross-validation cannot be performed on the target domain due to the absence of labels and must instead be performed on the training (source) data. In light of this, we adopt the prior knowledge approach proposed in [20].
The prior knowledge approach requires that side information about the similarities between the domains in the target domain is known by the target client. In freda, this approach involves the target client partitioning the indices $\{1,\cdots,l\}$ of all domains available in the target domain data $X^{T}$, creating two subsets $X^{t_{1}}$ and $X^{t_{2}}$. Afterwards, source clients federatively train separate weighted elastic nets, where during the training of each model the $\lambda$ value is selected by varying different possible values on a grid. Then, all these models trained with different values of $\lambda$ are sent to the target client, where the target client picks the model that leads to the lowest MAE on the target domains in $X^{t_{1}}$, assuming that the corresponding labels are available. Next, the target client fits a simple linear model for the relationship between domain similarity and the optimal $\lambda$ values obtained from the weighted elastic nets provided by the source clients. Using this model, the target client predicts the optimal $\lambda$ values for all domains in the remaining partition $X^{t_{2}}$. These predicted values are then sent to the source domain owners for the final phase.
5.3.2 Final Adaptive Model Training
The source clients federatively train the final weighted elastic net models by selecting the regularization parameter $\lambda$ based on the predicted values received from the target client in the previous step.
The model obtained at the end of this step is sent to the target client by the aggregator for the final prediction task on the target domain data.
@@ 6 Security Analysis
  3. 3.
    The communication channel between the aggregator and the clients is secure. It is not possible for other clients to perform man-in-the-middle attacks or eavesdrop on the exchanged messages.
@@ 6.1 Security of the Federated Feature Model Training
The training of feature models consists of two key steps (Section 5.1): federated hyper-parameter optimization and federated GPR training.
@@ 6.1.1 Security of the Federated Hyper-parameter Optimization
This step is performed collaboratively between the source clients and the aggregator. The local data of the source clients are never exposed directly to either the aggregator or other data owners during this process. The only information exchanged during this step includes the optimal values of the prior variance on the coefficients $\sigma_{p}^{2}$ and the variance of the additive noise $\sigma_{n}^{2}$ (from the closed-form solution in Eq. 2). These values are computed locally by each source client for each feature.
The use of secure aggregation, implemented via zero-sum masking as formally proven in [38], ensures that the semi-honest aggregator cannot infer individual hyper-parameter values for any source client. Instead, the aggregator only gains access to the aggregated sum of hyper-parameters for each feature, which contains no exploitable information about the local datasets.
@@ 6.1.2 Security of the Federated GPR Training
This step is performed collaboratively between all participants, which include the target client, source clients, and the aggregator.
@@
Aggregator’s Perspective.  The information disclosed to the aggregator is limited to the number of local data samples held by each source client, as indicated by the size of each sub-matrix (Section 5.1.2). This information is necessary for the correct partitioning of the intermediate mean matrix $K_{*}K^{-1}$, but it does not reveal any details about the content of the local data samples. Due to the masking process employed, the aggregator cannot infer any information regarding the local datasets of the source or target clients. The only data accessible to the aggregator is the resulting Gram matrix $G_{pq}$, which is computed from the masked matrices. The security of this masking process, implemented within the FLAKE framework, is formally proven in [37].
Source Clients’ Perspective.  The information available to each source client is limited to the respective sub-matrix they receive from the aggregator, which is derived from $K_{*}K^{-1}$. The aggregator masks the intermediate matrix product by multiplying it with a mask matrix $C$ (Section 5.1.2), resulting in the masked matrix $CK_{*}K^{-1}$. This masked matrix is then split into sub-matrices and sent to the respective source clients. Since each source client only has access to their own sub-matrix, they cannot infer any information about the overall matrix or the data of other clients without having access to the masking matrix $C$.
Target Client’s Perspective.  The information available to the target client includes the masked mean vectors sent by the source clients, the inverse of the masking matrix $C^{-1}$, and the predicted variance $K_{**}-K_{*}K^{-1}K_{*}^{\top}$ sent by the aggregator. To remove the mask on the mean vectors, the target client computes:
$K_{*}K^{-1}y=C^{-1}\sum_{i=1}^{N}(CK_{*}K^{-1})^{s_{i}}y^{s_{i}}.$
Since the individual vectors from the source clients are aggregated during this process, the target client cannot infer any meaningful information about the local datasets of the source clients. Regarding the predicted variance $K_{**}-K_{*}K^{-1}K_{*}^{\top}$, although the target client has access to $K_{**}$, it can compute $-K_{*}K^{-1}K_{*}^{\top}$. However, without access to the individual data matrices of the source clients, the target client cannot infer any meaningful information, as there are infinitely many possible source data matrices that would result in the same $K$ and $K_{*}$ [37].
6.2 Security of the Feature Weight Computation
@@
The local data of the target client is never directly exposed to either the aggregator or the source clients during this process. The only information shared with the aggregator and source clients consists of the feature weights for each tissue in the target domain. These feature weight vectors are aggregate values, computed by the target client by averaging the feature weights of the samples belonging to each tissue within the target domain data (Section 5.2). Consequently, it is not possible for the aggregator or the source clients to infer any meaningful information on the target data.
@@ 6.3 Security of the Optimal Lambda Prediction
Aggregator’s Perspective.  During the federated training process, the aggregator receives masked model updates from the source clients, which it aggregates to compute the coefficients of the global model for the next iteration. The use of secure aggregation, implemented via zero-sum masking as formally proven in [38], ensures that a semi-honest aggregator cannot infer individual updates from any source client.
@@
Target Client’s Perspective.  The target client receives only the coefficients of models trained using various regularization parameters. These model coefficients are aggregate values that do not reveal any specific information about the local data of any source client.
@@ 6.4 Security of the Final Adaptive Model Training
Aggregator’s Perspective.  The aggregator receives masked model parameters from the source clients, which it aggregates to compute the coefficients of the global model for the next iteration. The use of secure aggregation, implemented via zero-sum masking as formally proven in [38], ensures that a semi-honest aggregator cannot infer individual updates from any source client.
Target Client’s Perspective.  The target client receives only the coefficients of models trained using various regularization parameters. These model coefficients are aggregate values that do not reveal any specific information about the local data of any source client.
@@ 7 Conclusion
In this article, we introduce freda, the first method to enable privacy-preserving, federated unsupervised domain adaptation for regression tasks in high-dimensional, small-scale biological datasets. Unlike existing approaches that rely on deep learning or centralized training, freda allows multiple entities to collaboratively model complex feature relationships while maintaining complete data privacy. By leveraging a novel combination of randomized encoding and secure aggregation, our framework overcomes the fundamental challenge of training Gaussian Processes in federated settings by solving the need for centralized pairwise computations on data that cannot be directly shared.
Gaussian Processes are well-suited for modeling high-dimensional, small-sample datasets due to their probabilistic nature and ability to provide uncertainty estimates. However, their reliance on full data access has historically made them impractical for distributed learning scenarios. Freda is the first method to overcome this limitation, demonstrating that federated training of Gaussian Processes is both feasible and effective in real-world scenarios where privacy constraints and data heterogeneity present significant challenges.
To evaluate freda, we applied it to a benchmark task of age prediction from DNA methylation data, comparing its performance against the state-of-the-art centralized method wenda and the non-adaptive baseline en-ls. Our results show that freda achieves comparable performance to the state-of-the-art while ensuring complete privacy and enabling cross-institutional collaboration without direct data sharing. Our findings highlight the potential of freda as a generalizable framework for domain adaptation in high-dimensional biological datasets.
By introducing the first privacy-preserving, federated training framework which utilize Gaussian Processes in domain adaptation, this work provides a foundation for secure, data-efficient modeling in computational biology and other fields where data privacy, heterogeneity and sample volume are major concerns.